OpenStudy (anonymous):

A system of equations is shown below: Equation A: 3c = d − 8 Equation B: c = 4d + 8 Which of the following steps should be performed to eliminate variable d first?

OpenStudy (anonymous):

Equation B tells you what c equals so we sub in c for where it says 3c \[\huge ~\rm~3(4d+8)=d-8\]

OpenStudy (anonymous):

Distribute the 3 \[\huge ~\rm~3(4d+8)\]

OpenStudy (anonymous):

A system of equations is shown below: n = 3m + 7 n − 2m = 1 What is the solution, in the form (m, n), to the system of equations?

OpenStudy (anonymous):

It tells you what n equals 3m+7-2m=1 combine like terms to get the m value

OpenStudy (anonymous):

(-6,-11)?